1.Traveling 900km by rail costs Rs 280, What would be the fare for a journey of 360km when a person travel by the same class.
2. 3/4 of the salary of per month is Rs 600. What is the salary per month?

Answer by checkley79(3341)

900/280=360/X CROSS MULTIPLY.
900X280*360
900X=100,800
X=100,800/900
X=112 rS. WOULD BE THE COST.
########################################
3/4*X=600
3X/4=600
3X=4*600
3X=2,400
X=2,400/3
X=800 Rs. ANS.